A treasure map is drawn to a scale of 1 inch =5 miles. On the map ,the distance between Point A and X-marks-the-spot is 2 3/5. How many actual miles does this represent?
Answer by checkley79(3341)

(2 3/5)*5=10 15/5=10+3=13 miles ans.